A rectangular lawn has sides which measure x − 5 feet and 3x feet. The perimeter is 46 feet. Find the length of each side of the lawn.
Answer by ikleyn(52786)

.
The perimeter = (x-5) + 3x + (x-5) + 3x = 8x - 10 feet,


so your equation for x is


    8x - 10 = 46,   or

    8x = 46 + 10 = 56,   which implies

     x = 56/8 = 7 feet.


Then the sides of the lawn are  x-5 = 7-5 = 2 ft   and   3x = 3*7 = 21 feet.    ANSWER

Solved.

Check if you did not forget to post your "Thanks" to me . . .